# Graph Convolutions Enrich the Self-Attention in Transformers!

### v2

@@ Appendix C Implementation of GFSA
Appendix D Oversmoothing and Additional Visualizations
In Figure 2, we show the visualizations of oversmoothing characteristics in DeiT. We also provide visualizations in other domains. We show the filter response, cosine similarity, and singular value of BERT finetuned on STS-B dataset of GLUE tasks in Figure 4 and Graphormer finetuned on ZINC dataset in Figure 5.
To characterize self-attention, we first analyze the filter response of self-attention in the frequency domain. We follow the method used by Wang et al. (2022) for spectral visualization of the self-attention matrix. As shown in Figure 2 (a), Deit has a near-zero magnitude for the high frequencies, which is characteristic of a low-frequency filter and is likely to result in oversmoothing when applied multiple times.
We follow the calculation method of Guo et al. (2023) for cosine similarity. As shown in Figure 2 (b), the higher similarity as the layers of the model get deeper is related to the oversmoothing problem. To further analyze this issue, we also consider the dimensionality collapse in Transformer-based models. We plot the singular value distribution of the feature in the last block. As shown in Figure 2 (c), insignificant, near-zero values dominate the feature distribution. As layers get deeper, the similarity of features increases and dimensional collapse occurs. The oversmoothing problem is the same in BERT and Graphormer, as shown in Figure 4 and Figure 5.
(a) Filter reponse
(b) Cosine Similarity
(c) Singular Value
(a) Filter reponse
(b) Cosine Similarity
(c) Singular Value
Appendix E Natural Language Understanding
E.1 Detailed Experimental Settings
@@
CoLA
The Corpus of Linguistic Acceptability (Warstadt et al., 2019) consists of English acceptability judgments drawn from books and journal articles. The target task is a binary classification task, and each sentence is determined to be grammatically acceptable or not.
SST-2
The Stanford Sentiment Treebank (Socher et al., 2013) is a dataset in which each sentence is sourced from movie reviews and accompanied by human annotations of their sentiment. The target task is binary sentiment classification for a single sentence.
MRPC
The Microsoft Research Paraphrase Corpus (Dolan & Brockett, 2005) is a corpus of sentence pairs, which are automatically extracted from online news sources and annotated by humans. The target is to determine whether the sentences in the pair are semantically equivalent.
QQP
The Quora Question Pairs (Chen et al., 2018) dataset is a collection of question pairs from the community question-answering website Quora. The target is to determine whether the questions in the pair are semantically equivalent.
STS-B
The Semantic Textual Similarity Benchmark (Cer et al., 2017) is a collection of sentence pairs drawn from news headlines, video and image captions, and natural language inference data with human annotation. The target is a regression task to predict a similarity score from 0 to 5.
MNLI
The Multi-Genre Natural Language Inference Corpus (Williams et al., 2017) is a crowdsourced collection of sentence pairs with textual entailment annotations. Given a premise sentence and a hypothesis sentence, the task is to predict whether the premise entails the hypothesis (entailment), contradicts the hypothesis (contradiction), or neither (neutral). The standard test set consists of private labels from the authors and evaluates both the matched (in-domain) and mismatched (cross-domain) sections.
QNLI
The Stanford Question Answering (Wang et al., 2018) dataset is a question-answering dataset consisting of question-paragraph pairs, where one of the sentences in the paragraph contains the answer to the corresponding question written by an annotator. The task is to determine whether the context sentence contains the answer to the question.
RTE
The Recognizing Textual Entailment (Bentivogli et al., 2009) dataset comes from a series of annual textual entailment challenges. The target task is a binary entailment classification task.
@@
BERT (Devlin et al., 2019) consists with 12 layers, 12 heads, 768 hidden size, 512 maximum sequence length, and MLP dimension of 3072.
@@
ALBERT (Lan et al., 2019) consists with 12 layers, 12 heads, 768 hidden size, 512 maximum sequence length, 128 embedding dimension, and MLP dimension of 3072.
@@
RoBERTa (Liu et al., 2019) consists with 12 layers, 12 heads, 768 hidden size, 514 maximum sequence length, and MLP dimension of 3072.
@@
E.2 Sensitivity to $K$
In this section, we explore the influence of the polynomial order, denoted as $K$, in our GFSA, conducting experiments on BERTBASE finetuned with GLUE tasks. We search for values of $K$ from 2 to 10, and the results are presented in Figure 6. For each dataset, Optimal $K$ exists and the performance of models using GFSA is generally robust to changes in $K$.
(a) CoLA
(b) SST-2
(c) MRPC
(d) QQP
(e) STS-B
(f) MNLI-m
(g) MNLI-mm
(h) QNLI
(i) RTE
Figure 6: Sensitivity results on various $K$ with BERTBASE finetuned on GLUE tasks
Appendix F Causal Language Modeling
F.1 Detailed Experimental Settings
@@
PTB
Penn Treebank (Marcus et al., 1993) dataset is a collection of text documents that have been extensively annotated with linguistic information, primarily syntactic and grammatical structures.
WikiText
WikiText (Merity et al., 2016) dataset is a collection of over 100 million tokens extracted from the set of verified good and featured articles on Wikipedia. Compared to the preprocessed version of Penn Treebank (PTB), WikiText-2 is over 2 times larger and WikiText-103 is over 110 times larger.
@@
GPT-2 (Radford et al., 2019) is a Transformer pretrained on a very large corpus of English data in a self-supervised fashion without any human labelling on dataset. It automatically generate inputs and labels from those texts, and trained to guess the next word in sentences. For implementation, we adopt HuggingFace Framework  [^3] https://github.com/huggingface/transformers. For all experiments, GPT2 has 12 layers with 12 attention heads, 768 hidden size and 1024 maximum sequence length, resulting in a total of 117 million parameters.
@@ F.2 Sensitivity to $K$
Appendix G Image Classification
G.1 Detailed Experimental Settings
Our code is implemented based on the timm library (Wightman, 2019). In the case of our training recipe, it is the same as experimental setting of Wang et al. (2022) that follows the training recipes of Touvron et al. (2021a) and Touvron et al. (2021b). To apply our GFSA to existing base models such as DeiT, Cait, and Swin, we consider a range of $K$ between 2 and 5. For 12-layer DeiT, we follow the same hyperparameters from Wang et al. (2022). We set the dropout rate to 0 and 0.2 for 12-layer and 24-layer DeiT, respectively. For CaiT, we apply our GFSA on only to the patch embedding layer. All other hyper-parameters are kept consistent with the original papers of DeiT Touvron et al. (2021a), CaiT (Touvron et al., 2021b) and, Swin (Liu et al., 2021). All models are trained on NVIDIA RTX 3090 24GB.
G.2 FLOPs & Throughput
@@
G.3 Sensitivity to $K$
We also perform the sensitivity analysis for $K$. Tables 11 and 11 show the results of sensitivity analysis for DeiT-S and CaiT-S with GFSA plugged in. For 12-layer DeiT-S, GFSA performance of 81.12 is highest when $K=3$. When the GFSA has a $K$ of 2, the performance is worse than the original DeiT-S, but when the $K$ is 3 or higher, the performance is better than the original DeiT-S, and most surprisingly, the performance is better than the 24-layer DeiT-S.
@@ G.3 Sensitivity to $K$
Table 11: Sensitivity to $K$ for 24-layer CaiT-S + GFSA
G.4 Additional Experiments with Guo et al. (2023)’s setting
To make a fair comparison with ContraNorm (Guo et al., 2023), one of the related studies that mitigates oversmoothing, we run additional experiments to match their experimental setup.
Experimental Setting.
We follow the training recipe used by Guo et al. (2023), which is a slightly modified version of Touvron et al. (2021a)’s recipe. Guo et al. (2023) use AdamW optimizer with cosine learning rate decay. We select the DeiT-T and DeiT-S for ImageNet-1k. “T” and “S” denote tiny and small model sizes, respectively. For all experiments, the image size is set to be 224x224. We train each model for 300 epochs and the batch size is set to 1024. For ContraNorm, we train with their recommended hyperparameters. All models are trained on 4 GPUs and of NVIDIA RTX A6000 48GB.
Experimental Results.
@@
Appendix H Graph Classification
H.1 Detailed Experimental Settings
@@
The benchmark datasets we used are listed below.
ZINC
ZINC (Irwin et al., 2012) is the most popular real-world molecular dataset to predict graph property regression for constrained solubility, an important chemical property for designing generative GNNs for molecules. Uniform sampling is adopted for data splitting. We use a ZINC-subset of small-scale dataset.
PCQM4M
PCQM4M-LSC  (Hu et al., 2021) is 2D molecular graphs, which is one of the most practically relevant quantum chemical properties of molecule science. The task is to predict DFT(density functional theory)-calculated HOMO-LUMO energy gap of molecules given their graphs. PCQM4M-LSC is unprecedentedly large in scale comparing to other labeled graph-level prediction datasets, which contain more than 3.8M graphs. We use PCQM4M and PCQM4Mv2 large-scale datasets.
Graphormer.
Following Graphormer (Ying et al., 2021), we use Graphormer for PCQM4M and GraphormerSLIM for ZINC. Graphormer consists of 12 encoder layers, 80 encoder embedding dimension, and 768 MLP dimension. It employs 32 encoder heads and 24 hidden dimension for each head. GraphormerSLIM consists of 12 encoder layers, 80 encoder embedding dimension, and 80 MLP dimension. It employs 8 encoder heads and 10 hidden dimension for each head.
Training.
[…]
H.2 Experimental results with standard deviation
Table 14: Experimental results and number of parameters on ZINC
Method #Params MAE
Graphormer 500K 0.1240±0.006
Graphormer + GFSA 500K 0.1189±0.002
Method #Params PCQM4M PCQM4Mv2
Train Validate Train Validate
Graphormer 48.3M 0.0535±0.038 0.1286±0.016 0.0250±0.000 0.0862±0.000
Graphormer + GFSA 48.3M 0.0312±0.001 0.1193±0.000 0.0249±0.000 0.0860±0.000
We conducted experiments following the experimental environments of Graphormer Ying et al. (2021) using 4 different seeds. Due to space constraints, only the mean values are reported in Tables 6 and 6. In this section, we report the results with mean and standard deviations in Tables 14 and 15.
Appendix I Automatic Speech Recognition
I.1 Detailed Experimental Settings
Dataset.
We conduct experiments on the LibriSpeech  [^4] http://www.openslr.org/12 dataset (Panayotov et al., 2015), which consists of audio recordings paired with their transcriptions. The LibriSpeech dataset has approximately 1,000 hours of read English speech with a sampling rate of 16 kHz. We keep the original 16, [^0] Hz sampling rate and compute 80-dim log-Mel filterbanks for a 25ms sliding window, strided by 10ms. The filterbank features are then normalized to zero mean and unit variance per input sequence. For implementation, we follow the recipes of SpeechBrain (Ravanelli et al., 2021).
@@ I.1 Detailed Experimental Settings
We use a vanilla Transformer to apply our GFSA. For implementation, we use a SpeechBrain (Ravanelli et al., 2021) framework. The vanilla Transformer consists of i) 1D convolution to perform striding, ii) Transformer encoder with 12 layers, 4 heads, embedding dimension of 512, MLP dimension of 2048, and post-LayerNorm iii) decoder with 6 layers, 4 heads, embedding dimension of 512, MLP dimension of 2048, joint beamsearch, and iv) external Transformer language model with 12 layers, 12 heads, embedding dimension of 768, and MLP dimension of 3072.
@@
We use one of the SOTA models, Branchformer (Peng et al., 2022) to plug-in our GFSA. Branchformer has two parallel branches, one for capturing global interactions using attention and the other for more localized context using convolutional gating MLP. The Branchformer architecture for speech recognition consists of i) 1D convolution to perform striding, ii) Branchformer encoder with 18 layers, 8 heads, embedding dimension of 512, and MLP dimension of 3072, iii) decoder with 6 layers, 8 heads, embedding dimension of 512, a convolutional spatial gating unit (CSGU) dimension of 3072, joint beamsearch, and iv) external Transformer language model with 12 layers, 12 heads, embedding dimension of 768, and MLP dimension of 3072.
@@
We follow a training recipe from SpeechBrain (Ravanelli et al., 2021). The standard LibriSpeech validation sets (dev-clean and dev-other) are used to tune all parameters and select the best models. Test sets (test-clean and test-other) are used only to report final WER performance. We train the pure Transformer for 100 epochs and the Branchformer for 120 epochs with a batch size of 16. We use a data augmentation method on all models using SpecAugment (Park et al., 2019). SpecAugment applies time and frequency masking as well as time warping to the input spectrum. For Branchformer, we use AdamW (Loshchilov & Hutter, 2017) optimizer with 0.9 and 0.98 coefficients for computing running averages of gradient and its square. The learning rate and weight decay in all models are 0.0008 and 0.01, respectively. We use a connectionist temporal classification (CTC) loss Graves & Jaitly (2014); Zhang et al. (2016). We also apply dropout with probability 0.1 and label smoothing with weight 0.1 to mitigate overfitting. We fix the random seed as 74443 on all experiments. All models are trained on 1 GPU and of NVIDIA RTX A6000 48GB.
@@
I.2 Training Curve
We compare the training and validation curves for LibriSpeech 100h in Figure 7. The training loss curve of GFSA is lower than the pure Transformer. GFSA stabilizes the loss curve of pure Transformer slightly earlier.
@@ I.2 Training Curve
Appendix J Code Defect Detection
J.1 Detailed Experimental Settings
@@
We use Devign dataset provided by (Zhou et al., 2019), which is a binary classification task to evaluate whether a C language code is vulnerable to software systems or not.
@@
We build our experiments on top of the open-sourced code  [^5] https://github.com/salesforce/CodeT5 and recipes provided by Wang et al. (2021b).
@@
RoBERTa (Liu et al., 2019) is an encoder-only model trained with masked language modeling on code. All hyperparameters are consistent with the training method in the source code of Wang et al. (2021b).
@@
PLBART (Ahmad et al., 2021) is an encoder-decoder model based on BART (Lewis et al., 2020) architecture. PLBART can support understanding and generation tasks. All hyperparameters are consistent with the training method in the source code of Wang et al. (2021b).
@@
CodeBERT (Feng et al., 2020) is a model trained on masked language modeling and replaced token detection. CodeBERT is a bimodal pretrained model based on Transformer with 12 layers for programming language and natural language. All hyperparameters are consistent with the training method in the source code of Wang et al. (2021b).
@@
CodeT5 is an encoder-decoder framework with the same architecture as T5 (Raffel et al., 2020). It aims to derive generic representations for programming language and natural language via pre-training on unlabeled source code. CodeT5-small has 6 encoder layers, 6 decoder layers, 8 attention heads, 512 dimensional hidden states, and 60M parameters. The other models have 12 encoder layers, 12 decoder layers, 12 attention heads, 768 dimensional hidden states, and 220M parameters. All hyperparameters are consistent with the training method in the source code of Wang et al. (2021b).

### v5

@@ Appendix I Implementation of GFSA
Appendix J Comparison with Actual and Approximated High-order Terms
To compare the impact of the actual $\bar{\bm{A}}^{K}$ and the approximated $\bar{\bm{A}}^{K}$ in terms of accuracy, we experimented with BERT on GLUE and the results are summarized in Table 9. BERTBASE+$\bar{\bm{A}}^{K}$ denotes using the exactly calculated $\bar{\bm{A}}^{K}$ instead of the approximated $\bar{\bm{A}}^{K}$.
Table 9: Comparison of performance using the exactly calculated $\bar{\bm{A}}^{K}$ vs. the approximated $\bar{\bm{A}}^{K}$ for GLUE tasks
Datasets #Params CoLA SST-2 MRPC QQP STS-B MNLI-m/mm QNLI RTE Avg
BERTBASE [16] 110M 56.79 93.81 88.70 88.32 88.16 84.96/84.15 91.63 66.06 82.51
+ GFSA (approximated $\bar{\bm{A}}^{K}$) 110M 59.56 94.15 90.60 88.46 88.33 85.12/85.06 91.95 68.95 83.58
+ GFSA (actual $\bar{\bm{A}}^{K}$) 110M 59.85 94.27 89.80 88.43 88.32 84.95/84.89 91.76 68.23 83.39
Appendix K Natural Language Understanding
K.1 Detailed Experimental Settings
We integrate GFSA into 3 pre-trained large language models: BERT, ALBERT, and RoBERTa. We evaluate them on the GLUE benchmark, which includes 3 categories of natural language understanding tasks: i) single-sentence tasks CoLA and SST-2; ii) similarity and paraphrasing tasks MRPC, QQP, and STS-B; iii) natural language inference tasks MNLI, QNLI, and RTE. For MNLI task, we experiment on both the matched (MNLI-m) and mismatched (MNLI-mm) versions. Following Devlin et al. [16], we report Matthews correlation for CoLA, F1 scores for QQP and MRPC, Spearman correlations for STS-B, and accuracy scores for the other tasks. For each task, we select the best hyperparameters for GFSA, and the other hyperparameters are fixed. We compare our GFSA with ContraNorm [28], one of the related methods that address oversmoothing. We finetune ContraNorm with the recommended hyperparameters in Guo et al. [28]. We initialize with a pre-trained language model and finetune with added GFSA for 5 epochs.
@@
  • CoLA. The Corpus of Linguistic Acceptability [85] consists of English acceptability judgments drawn from books and journal articles. The target task is a binary classification task, and each sentence is determined to be grammatically acceptable or not.
  • SST-2. The Stanford Sentiment Treebank [73] is a dataset in which each sentence is sourced from movie reviews and accompanied by human annotations of their sentiment. The target task is to classify binary sentiments for a single sentence.
  • MRPC. The Microsoft Research Paraphrase Corpus [17] is a corpus of sentence pairs, which are automatically extracted from online news sources and annotated by humans. The target is to determine whether the sentences in the pair are semantically equivalent.
  • QQP. The Quora Question Pairs [11] dataset is a collection of question pairs from the community question-answering website Quora. The target is to determine whether the questions in the pair are semantically equivalent.
  • STS-B. The Semantic Textual Similarity Benchmark [8] is a collection of sentence pairs drawn from news headlines, video and image captions, and natural language inference data with human annotation. The target is a regression task to predict a similarity score from 0 to 5.
  • MNLI. The Multi-Genre Natural Language Inference Corpus [87] is a crowdsourced collection of sentence pairs with textual entailment annotations. Given a premise sentence and a hypothesis sentence, the task is to predict whether the premise entails the hypothesis (entailment), contradicts the hypothesis (contradiction), or neither (neutral). The standard test set consists of private labels from the authors and evaluates both the matched (in-domain) and mismatched (cross-domain) sections.
  • QNLI. The Stanford Question Answering [82] dataset is a question-answering dataset consisting of question-paragraph pairs, where one of the sentences in the paragraph contains the answer to the corresponding question written by an annotator. The task is to determine whether the context sentence contains the answer to the question.
  • RTE. The Recognizing Textual Entailment [5] dataset comes from a series of annual textual entailment challenges. The target task is a binary entailment classification task.
@@
BERT [16] consists with 12 layers, 12 heads, 768 hidden size, 512 maximum sequence length, and MLP dimension of 3072.
@@
ALBERT [40] consists of 12 layers, 12 heads, 768 hidden dimensions, 512 maximum sequence length, 128 embedding dimensions, and MLP dimension of 3072.
@@
RoBERTa [44] consists of 12 layers, 12 heads, 768 hidden size, 514 maximum sequence length, and MLP dimension of 3072.
@@
K.2 Sensitivity to $K$
In this section, we explore the influence of the polynomial order, denoted as $K$, in our GFSA, conducting experiments on BERTBASE finetuned with GLUE tasks. We search for values of $K$ from 2 to 10, and the results are presented in Table 10. For each dataset, there is an optimal $K$ and the performance of models using GFSA is generally robust to changes in $K$.
Table 10: Sensitivity results on various $K$ with BERTBASE finetuned on GLUE tasks
$K$ CoLA SST2 MRPC QQP STSB MNLI-m MNLI-mm QNLI RTE
2 57.83 94.15 90.60 88.41 88.27 84.96 84.90 91.74 68.23
3 58.56 93.46 89.77 88.41 88.33 85.08 84.75 91.78 68.59
4 59.56 93.46 89.77 88.45 88.29 85.06 85.06 91.76 68.59
5 58.10 93.58 90.07 88.39 88.29 84.87 84.99 91.85 68.95
6 59.40 93.58 90.40 88.29 88.27 84.93 84.97 91.69 68.23
7 59.12 94.04 90.48 88.43 88.26 85.12 84.94 91.82 68.94
8 58.58 93.69 90.12 88.46 88.24 84.92 84.81 91.95 68.95
9 58.88 93.46 89.54 88.41 88.26 85.06 91.67 67.87 85.04
10 59.31 93.35 89.98 88.41 88.30 84.84 91.73 68.59 84.93
Appendix L Causal Language Modeling
L.1 Detailed Experimental Settings
@@
  • PTB. Penn Treebank [47] dataset is a collection of text documents that have been extensively annotated with linguistic information, primarily syntactic and grammatical structures.
  • WikiText. WikiText [50] dataset is a collection of over 100 million tokens extracted from the set of verified good and featured articles on Wikipedia. Compared to the preprocessed version of Penn Treebank (PTB), WikiText-2 is over 2 times larger and WikiText-103 is over 110 times larger.
@@
GPT2 [61] is a Transformer pretrained on a very large corpus of English data in a self-supervised fashion without any human labelling on dataset. It automatically generate inputs and labels from those texts, and trained to guess the next word in sentences. For implementation, we adopt HuggingFace Framework  [^4] https://github.com/huggingface/transformers. For all experiments, GPT2 has 12 layers with 12 attention heads, 768 hidden size and 1024 maximum sequence length, resulting in a total of 117 million parameters.
@@
Appendix M Image Classification
M.1 Detailed Experimental Settings
Our code is implemented based on the timm library [86]. In the case of our training recipe, it is the same as experimental setting of Wang et al. [80] that follows the training recipes of Touvron et al. [74] and Touvron et al. [75]. To apply our GFSA to existing base models such as DeiT, Cait, and Swin, we consider a range of $K$ between 2 and 5. For 12-layer DeiT, we follow the same hyperparameters from Wang et al. [80]. We set the dropout rate to 0 and 0.2 for 12-layer and 24-layer DeiT, respectively. For CaiT, we apply our GFSA on only to the patch embedding layer. All other hyper-parameters are kept consistent with the original papers of DeiT [74], CaiT [75] and, Swin [45]. All models are trained on NVIDIA RTX 3090 24GB.
M.2 FLOPs & Throughput
@@
M.3 Sensitivity to $K$
We also perform the sensitivity analysis for $K$. Tables 14 and 14 show the results of sensitivity analysis for DeiT-S and CaiT-S with GFSA plugged in. For 12-layer DeiT-S, GFSA performance of 81.12 is highest when $K=3$. When the GFSA has a $K$ of 2, the performance is worse than the original DeiT-S, but when the $K$ is 3 or higher, the performance is better than the original DeiT-S, and most surprisingly, the performance is better than the 24-layer DeiT-S.
@@
M.4 Full Experimental Results
In Table 15, we consider all three classes, CNN only, CNN + Transformer, and pure Transformer, to compare more different models than in Table 3. In particular, in the Transformer category, we only test with lightweight models with similar number of parameters, such as ViT-S and DeiT-S. Compared to existing techniques, the improvements by GFSA already surpasses LayerScale (0.7%) [75], LateInsertion (0.6%) [75], and HAT [4] (1.38%).
Category Method Input Size #Layers #Params Top-1 Acc
CNN ResNet-152 [29] 224 152 230M 78.1
DenseNet-201 [34] 224 201 77M 77.6
CNN + CVT-21 [88] 224 21 32M 82.5
Transformer Refiner [102] 224 16 86M 81.2
Transformer ViT-S/16 [19] 224 12 49M 78.1
ViT-B/16 [19] 224 12 86M 79.8
DeiT-S [74] 224 12 22M 79.8
DeiT-S + LayerScale [75] 224 12 22M 80.5
DeiT-S + LateInsertion [75] 224 12 22M 80.5
DeiT-S + ClassAttention [75] 224 12 22M 80.6
DeiT-S + AttnScale [80] 224 12 22M 80.7
DeiT-S + FeatScale [80] 224 12 22M 80.9
DeiT-S + HAT [4] 224 12 22M 80.9
DeiT-S + Diverse [10] 224 12 22M 80.6
DeiT-S + ContraNorm [28] 224 12 22M 80.4
Swin-S [45] 224 12 50M 82.9
T2T-ViT-24 [96] 224 24 64M 82.3
DeepViT-24B [101] 224 24 36M 80.1
DeiT-S [74] 224 24 43M 80.5
CaiT-S [75] 224 24 47M 82.6
DeiT-S + DiversePatch [25] 224 24 44M 82.2
DeiT-S + LayerScale [75] 224 24 44M 82.4
DeiT-S + AttnScale [80] 224 24 44M 81.1
DeiT-S + FeatScale [80] 224 24 44M 81.3
DeiT-S + ContraNorm [28] 224 24 43M 80.7
GFSA DeiT-S + GFSA 224 12 22M 81.1 ($\uparrow$ 1.3)
DeiT-S + GFSA 224 24 43M 81.5 ($\uparrow$ 1.0)
CaiT-S + GFSA 224 24 47M 82.8 ($\uparrow$ 0.2)
Swin-S + GFSA 224 12 50M 83.0 ($\uparrow$ 0.1)
M.5 Additional Comparison with SOTA Models
Our main experiment aims to determine whether introducing the GFSA layer would help improve performance in a base model, such as DeiT. We also compare our method with the recent models: SpectFormer [58], SVT [57], NeuTRENO [52], FNet [42], and GFNet [64]. We use a 12-layer setup to ensure a fair comparison in Table 16.
GFNet [64] can reduce the number of parameters, but there is a performance penalty. However, the performance improvement of DeiT-S+GFSA is relatively greater than DeiT-S compared to other models. SpectFormer [58] and SVT [57] have advantages in calculation amount and model complexity, and performance is improved over DeiT-S, but Top-1 and Top-5 accuracies are lower than those using GFSA. Additionally, NeuTRENO [52] also improves as much as GFSA compared to DeiT-S, but GFSA still has higher Top-1 accuracy.
Table 16: Compared with state-of-the-art models on ImageNet-1k
Method Input Size #Layers #Params Top-1 Acc Top-5 Acc
DeiT-S 224 12 22M 79.8 95.0
Fnet-XS [42] 224 12 20M 71.2 -
GFNet-XS [64] 224 12 16M 78.6 94.2
SpectFormer-XS [58] 224 12 20M 80.2 94.7
SVT-XS [57] 224 12 20M 79.9 94.5
DeiT-S + NeuTRENO [52] 224 12 20M 80.7 95.4
DeiT-S + GFSA 224 12 22M 81.1 95.4
M.6 Additional Experiments with Guo et al. [28]’s setting
To make a fair comparison with ContraNorm [28], one of the related studies that mitigates oversmoothing, we run additional experiments to match their experimental setup.
Setting.
We follow the training recipe used by Guo et al. [28], which is a slightly modified version of Touvron et al. [74]’s recipe. Guo et al. [28] use AdamW optimizer with cosine learning rate decay. We select the DeiT-T and DeiT-S for ImageNet-1k. “T” and “S” denote tiny and small model sizes, respectively. For all experiments, the image size is set to be 224x224. We train each model for 300 epochs and the batch size is set to 1024. For ContraNorm, we train with their recommended hyperparameters. All models are trained on 4 GPUs and of NVIDIA RTX A6000 48GB.
Results.
@@
Appendix N Automatic Speech Recognition
N.1 Detailed Experimental Settings
@@
We conduct experiments on the LibriSpeech  [^5] http://www.openslr.org/12 dataset [55], which consists of audio recordings paired with their transcriptions. The LibriSpeech dataset has approximately 1,000 hours of read English speech with a sampling rate of 16 kHz. We keep the original 16, [^0] Hz sampling rate and compute 80-dim log-Mel filterbanks for a 25ms sliding window, strided by 10ms. The filterbank features are then normalized to zero mean and unit variance per input sequence. For implementation, we follow the recipes of SpeechBrain [65].
@@
We use a vanilla Transformer to apply our GFSA. For implementation, we use a SpeechBrain [65] framework. The vanilla Transformer consists of i) 1D convolution to perform striding, ii) Transformer encoder with 12 layers, 4 heads, embedding dimension of 512, MLP dimension of 2048, and post-LayerNorm iii) decoder with 6 layers, 4 heads, embedding dimension of 512, MLP dimension of 2048, joint beamsearch, and iv) external Transformer language model with 12 layers, 12 heads, embedding dimension of 768, and MLP dimension of 3072.
@@
We use one of the SOTA models, Branchformer [59] to plug-in our GFSA. Branchformer has two parallel branches, one for capturing global interactions using attention and the other for more localized context using convolutional gating MLP. The Branchformer architecture for speech recognition consists of i) 1D convolution to perform striding, ii) Branchformer encoder with 18 layers, 8 heads, embedding dimension of 512, and MLP dimension of 3072, iii) decoder with 6 layers, 8 heads, embedding dimension of 512, a convolutional spatial gating unit (CSGU) dimension of 3072, joint beamsearch, and iv) external Transformer language model with 12 layers, 12 heads, embedding dimension of 768, and MLP dimension of 3072.
@@
We follow a training recipe from SpeechBrain [65]. The standard LibriSpeech validation sets (dev-clean and dev-other) are used to tune all parameters and select the best models. Test sets (test-clean and test-other) are used only to report final WER performance. We train the pure Transformer for 100 epochs and the Branchformer for 120 epochs with a batch size of 16. We use a data augmentation method on all models using SpecAugment [56]. SpecAugment applies time and frequency masking as well as time warping to the input spectrum. For Branchformer, we use AdamW [46] optimizer with 0.9 and 0.98 coefficients for computing running averages of gradient and its square. The learning rate and weight decay in all models are 0.0008 and 0.01, respectively. We use a connectionist temporal classification (CTC) loss [26, 100]. We also apply dropout with probability 0.1 and label smoothing with weight 0.1 to mitigate overfitting. We fix the random seed as 74443 on all experiments. All models are trained on 1 GPU and of NVIDIA RTX A6000 48GB.
@@
N.2 Training Curve
We compare the training and validation curves for LibriSpeech 100h in Fig. 8. The training loss curve of GFSA is lower than the pure Transformer. GFSA stabilizes the loss curve of pure Transformer slightly earlier.
@@
Appendix O Graph-level Tasks
O.1 Detailed Experimental Settings
Experimental settings for Graphormer.
We describe benchmark datasets and Graphormer as the backbone model we used. ZINC [35] is the most popular real-world molecular dataset to predict graph property regression for constrained solubility, an important chemical property for designing generative GNNs for molecules. Uniform sampling is adopted for data splitting. We use a ZINC-subset of small-scale dataset. PCQM4M-LSC  [33] is 2D molecular graphs, which is one of the most practically relevant quantum chemical properties of molecule science. The task is to predict density functional theory (DFT)-calculated HOMO-LUMO energy gap of molecules given their graphs. PCQM4M-LSC is unprecedentedly large in scale comparing to other labeled graph-level prediction datasets, which contain more than 3.8M graphs. We use PCQM4M and PCQM4Mv2 large-scale datasets.
[…]
Experimental settings for GPS and Graph-ViT.
We use various benchmark datasets to experiment with our GFSA on GPS [63] and Graph-ViT [30]: Peptide-func and Peptide-struct from Long-Range Graph Benchmark (LRGB) [21], MNIST, CIFAR10, and ZINC from Benchmarking GNNs [22], and Moltox21 and Molhiv from OGB [32]. We fix all hyperparameters of GPS and Graph-ViT as recommended in their paper to ensure a fair comparison. To plug GFSA into GPS, we replace the self-attention module of GPS with GFSA. For Graph-ViT, we apply GFSA instead of the Hadamard self-attention mechanism and compare it with this self-attention method. We conduct experiments on GPS and Graph-ViT in their open code frameworks for a fair comparison:
O.2 Experimental Results with Standard Deviation
We conduct experiments following the experimental environments of Graphormer [94] using 4 different seeds. Due to space constraints, only the mean values are reported in Tables 5 and 5. In Tables 20 and 21, we report the results with mean and standard deviations.
Table 20: Experimental results and number of parameters on ZINC
Method #Params MAE
Graphormer 500K 0.1240$\pm$0.006
Graphormer + GFSA 500K 0.1189$\pm$0.002
Method #Params PCQM4M PCQM4Mv2
Train Validate Train Validate
Graphormer 48.3M 0.0535$\pm$0.038 0.1286$\pm$0.016 0.0250$\pm$0.000 0.0862$\pm$0.000
Graphormer + GFSA 48.3M 0.0312$\pm$0.001 0.1193$\pm$0.000 0.0249$\pm$0.000 0.0860$\pm$0.000
Appendix P Code Defect Detection
P.1 Detailed Experimental Settings
@@
We use Devign dataset provided by [105], which is a binary classification task to evaluate whether a C language code is vulnerable to software systems or not.
@@
We build our experiments on top of the open-sourced code  [^6] https://github.com/salesforce/CodeT5 and recipes provided by Wang et al. [84].
@@
RoBERTa [44] is an encoder-only model trained with masked language modeling on code. All hyperparameters are consistent with the training method in the source code of Wang et al. [84].
@@
PLBART [2] is an encoder-decoder model based on BART [43] architecture. PLBART can support understanding and generation tasks. All hyperparameters are consistent with the training method in the source code of Wang et al. [84].
@@
CodeBERT [23] is a model trained on masked language modeling and replaced token detection. CodeBERT is a bimodal pretrained model based on Transformer with 12 layers for programming language and natural language. All hyperparameters are consistent with the training method in the source code of Wang et al. [84].
@@
CodeT5 is an encoder-decoder framework with the same architecture as T5 [62]. It aims to derive generic representations for programming language and natural language via pre-training on unlabeled source code. CodeT5-small has 6 encoder layers, 6 decoder layers, 8 attention heads, 512 dimensional hidden states, and 60M parameters. The other models have 12 encoder layers, 12 decoder layers, 12 attention heads, 768 dimensional hidden states, and 220M parameters. All hyperparameters are consistent with the training method in the source code of Wang et al. [84].